A hospital needs to dilutea 60% boric acid solution to a 10% solution. If it needs 40 liters of the 10% solution, how much of the 60% solution and how much water should it use?
Answer by mananth(16949)

60% boric acid solution x liters
water 40-x liters
10% solution. 40 liters
..
0.6x +0 = 0.1*40
0.6x=4
divide by 0.6
x= 4/0.6
x= 6.66 liters
40-6.66 = 33.34 liters of water